A ball of mass M moving with speed u colloid with the floor at an angle Alpha with the floor .  if it rebounds at the same angle Alpha with the same speed. What is the change in momentum?

Grade:11

1 Answers

Khimraj
5 years ago
change in momentum = final momentum – initial momentum
 = M*u*sin\alpha – (-M*u*sin\alpha) = 2*M*u*sin\alpha
Hope it clears..........................................